OpenStudy (anonymous):

Find ker (T) a) T: R^2-->R^2; where T(x,y) = (y,x) b) T : R^2-->T^2; where T(x,y)= (0, 2x+3y) c) T: R^2---> R^3; where T(x,y)= (x-y,y-x,2x-2y) Please, help

OpenStudy (amistre64):

set it up as a solution to \(c_1v_1+c_2v_2=0\)

OpenStudy (amistre64):

where v1 and v2 .. vn are the vectors presented

OpenStudy (anonymous):

more, please, still not get

OpenStudy (amistre64):

T(x,y) = (y,x) is notation for a setup such that 0x + 1y = 0 1x + 0y = 0

OpenStudy (anonymous):

got the step

OpenStudy (amistre64):

T(x,y)= (0, 2x+3y) is notation for 0x + 0y = 0 2x + 3y = 0

OpenStudy (anonymous):

got it, too. but , please, each case at a time, don't mix them up

OpenStudy (amistre64):

T(x,y)= (x-y,y-x,2x-2y) 1x- 1y =0 -1x +1y =0 2x - 2y =0 the setup is then augmented to determine the free variables

OpenStudy (amistre64):

rref each coefficient matrix

OpenStudy (anonymous):

ok, then, teh rref tell me what? for case a) I got (0,0) b) I got (1,-3/2) so??

OpenStudy (amistre64):

the rref tells us the free variables and the ... not free? .... variables. forgot the names of the others :) so for the first one: 0 1 1 1 1 1 1 0 1 0 1 0 0 -1 0 1 has no free variables so the basis of the kernel is just the 0 vector

OpenStudy (anonymous):

got it

OpenStudy (amistre64):

for the second one 0 0 2 3 1 3/2 2 3 0 0 0 0 1 free variable such that x = -3/2 y y = 1 y the basis of our kernel is the vector <-3/2,1> , or to proper it up ... <3,-2>

OpenStudy (anonymous):

ok, got it

OpenStudy (amistre64):

1 -1 -1 1 2 -2 1 -1 0 0 0 0 in this case we have one free variable again, such that: x = 1y y = 1y z = 0y

OpenStudy (anonymous):

stuck here, 1 equation, 3 variables, we must have at least 2 free variables???

OpenStudy (amistre64):

what is the dim of the 3rd setup?

OpenStudy (anonymous):

1

OpenStudy (amistre64):

might have forgotten a specific along the way; the basis of the kernel on the last one is simply (1,1)

OpenStudy (anonymous):

so, I just take rref and consider how many free variables it has, and base on the result to conclude about the ker(T)?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

what if the matrix give out linear dependent . I mean 1 free variable and 2 independent . how to conclude about ker(T)?

OpenStudy (amistre64):

A = 1 -1 -1 1 2 -2 Ax = 0 when x = <1,1>

OpenStudy (amistre64):

define the 2 nonfree variables in terms of the free variable.

OpenStudy (anonymous):

the next step is " base on the ker(T) consider whether it is one to one or not." so I need how to find it, and how to consider ker(t) =0 or not

OpenStudy (amistre64):

cant say a recall the details of that

OpenStudy (anonymous):

in case a) matrix gives out linear independent, as you state, its ker (T) = (0,0) . got it

OpenStudy (anonymous):

case b) matrix gives out linear dependent, since it has free variable there, ker (T) = (-3,2) got it, too

OpenStudy (anonymous):

stuck at case c)

OpenStudy (anonymous):

hey, explanation please

OpenStudy (amistre64):

A linear transformation T is one-to-one ⇔ kernelT = {0} .... from the internet

OpenStudy (anonymous):

from my book, too

OpenStudy (amistre64):

if a system has a unique solution for Ax=b, then it is one to one

OpenStudy (amistre64):

if Ax = 0 has only the trivial solution, it is one to one

OpenStudy (anonymous):

I know, but the problem asks for find ker(T) and then consider.... Cannot come up by another way

OpenStudy (amistre64):

another property is such that T:R^n -> R^m is 1-1 if m>=n

OpenStudy (anonymous):

you mean we can conclude right there by the last property? it's not one to one.

OpenStudy (amistre64):

http://algebra.math.ust.hk/matrix_linear_trans/04_oneone/lecture2.shtml according to this, yes

OpenStudy (anonymous):

got it, thanks a lot, friend.

OpenStudy (amistre64):

hope it helps

OpenStudy (anonymous):

thanks for the link, I read it now